OpenStudy (gamereternity):

Apples sell for $1.90 per pound, and bananas sell for $0.75 per pound. Troy bought some apples and some bananas. Together they weighed 3.8 pounds, and cost $5.84. Let p represent the number of pounds of apples Troy purchased. Which equation represents the problem described here? A. 1.15p + 2.85 = 3.8<=my answer B. –1.15p + 2.85 = 5.84 C. 1.9p + 0.75 = 3.8 D. 1.15p + 2.85 = 5.84

OpenStudy (gamereternity):

@jim_thompson5910

OpenStudy (gamereternity):

@Mehek14

jimthompson5910 (jim_thompson5910):

`Let p represent the number of pounds of apples Troy purchased. ` p = the number of pounds of apples q = the number of pounds of bananas p+q = 3.8 since `Together they weighed 3.8 pounds` solve for q p+q = 3.8 p+q-p = 3.8-p q = 3.8-p -------------------------------------- `Apples sell for $1.90 per pound` buying p pounds of apples costs a total of 1.90p dollars `bananas sell for $0.75 per pound` buying q pounds of bananas costs a total of 0.75q dollars We have 1.90p+0.75q as the total cost in dollars. The total cost is known to be 5.84, so, 1.90p+0.75q = 5.48 Now plug in q = 3.8-p 1.90p+0.75q = 5.48 1.90p+0.75(3.8-p) = 5.48 1.90p+0.75(3.8)+0.75(-p) = 5.48 1.90p+2.85-0.75p = 5.48 1.15p+2.85 = 5.48 So you are close, but not quite there.

OpenStudy (gamereternity):

so b

jimthompson5910 (jim_thompson5910):

look more carefully at what I wrote

OpenStudy (gamereternity):

d?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (gamereternity):

ty